# Quantum thermodynamic derivation of the energy resolution limit in magnetometry

### v1

@@ Abstract
The atom’s magnetic moment is $\mu=\mu_{B}/q$, where $\mu_{B}$ is the Bohr magneton and $q=[(S(S+1)+I(I+1)]/S(S+1)$ is the nuclear slowing down factor with $S=1/2$ and $I$ the nuclear spin [52]. Using the SD cross sections listed in [52] and the respective relative velocities $\overline{v}$ we find for a number density $n=10^{14}~{}{\rm cm^{-3}}$ and volume $V=10~{}{\rm cm^{3}}$ that $\delta B=2\times 10^{-17}~{}{\rm T}$ for ${}^{41}{\rm K}$, $\delta B=10^{-16}~{}{\rm T}$ for ${}^{87}{\rm Rb}$, and $\delta B=10^{-14}~{}{\rm T}$ for ${}^{133}{\rm Cs}$. Using instead the parameters of [57], where the authors used a cesium vapor of volume $V=1~{}{\rm cm^{3}}$ and number density $2\times 10^{13}~{}{\rm cm^{-3}}$, we find again $\delta B\approx 10^{-14}~{}{\rm T}$. Since this noise is distributed within the bandwidth $1/\tau$, the corresponding spectral density is $10~{}{\rm pG/\sqrt{Hz}}$, whereas the authors measured a noise level of $400~{}{\rm pG/\sqrt{Hz}}$, and under the premise of a theoretical optimization project a noise level of $2~{}{\rm pG/\sqrt{Hz}}$. Using (3) we can obtain the ERL which reads
@@
Towards an experimental verification of the noise $\delta B$, we propose the use of a miniaturized cesium cell of volume e.g. $1~{}{\rm mm^{3}}$, in order to boost the previous estimates of $\delta B$ by a factor of $10^{3}$. The cell should be surrounded by a superconducting flux transformer in order to alleviate Johnson noise that would dominate the signal induced by the changing flux produced by $\delta B$ in an ordinary coil.
@@
There is yet another interpretation for the factor $\kappa$ in (5), interesting in its own right. We can consider the vacuum permeability $\mu_{0}$ is replaced by $\kappa\mu_{0}$, as if the spins have a tendency to align, and $\kappa$ represents the relative permeability constant. But indeed, binary spin-dependent collisions do have the tendency to correlate atoms, as was shown in [53] for the case of spin-exchange collisions. As mentioned before, spin-destruction collisions are described by a spin phase change $\phi$, which is $\phi\ll 1$. In [53] it was shown that the negativity of the two-atom spin state scales as $\sin\phi_{\rm se}^{2}$, where $\phi_{\rm se}\approx 1$ is the spin-exchange phase change. Similarly, here we can state that the entanglement between two colliding atoms will scale like $\sin\phi^{2}\approx\phi^{2}$. Hence any two colliding partners will share a small correlation. The upside is that this correlation can be shared by many more atoms. Imagine a ”bath” atom experiencing consecutive collisions with many ”system” atoms. These will all be slightly correlated, and will reside in a ”correlation volume” $V_{c}$. Any given spin within this volume will experience a magnetic field $\kappa\mu_{0}\mu/V_{c}$ and energy $\epsilon=\kappa\mu_{0}\mu^{2}/V_{c}$. This energy reorients the spin in a time $\hbar/\epsilon$. For the correlation to be maintained, this time should be equal or larger than $V_{c}/\sigma_{\rm sd}\overline{v}$, which is the time required for the bath atom to relax/correlate the volume $V_{c}$. From this requirement we obtain again $\kappa=\hbar\sigma_{\rm sd}\overline{v}/\mu_{0}\mu^{2}$. The correlation volume dropped out of this calculation, but we can still estimate it. We can write $\kappa=N_{c}\phi^{2}$, and $N_{c}=nV_{c}$. The quantity $\phi^{2}$ is related to the SD relaxation time and the collision rate $1/T$, i.e. it is $1/\tau=\phi^{2}/T$ as discussed in [56]. The collision rate is $1/T=\overline{v}/n^{-1/3}$. Putting everything together we find $V_{c}=\hbar\overline{v}/\mu_{0}\mu^{2}n^{5/3}$. For a number density of $n=10^{14}~{}{\rm cm}^{-3}$ it is $V_{c}\approx 0.01~{}{\rm mm}^{3}$, containing $N_{c}=10^{9}$ atoms.
@@
The physical picture that emerges is that both a polarized vapor being relaxed, as well as an unpolarized vapor, will behave like a “squashy” spin medium exhibiting spin fluctuations, split in a number of “domains” not unlike ferromagnetic systems. Such domains exhibit intra-domain correlations, but not inter-domain correlation. The vapor as a whole continuously exchanges energy with its self-field, setting an unavoidable noise level for measuring externally applied magnetic fields.

### v2

@@ Abstract
The atom’s magnetic moment is $\mu=\mu_{B}/q$, where $\mu_{B}$ is the Bohr magneton and $q=[(S(S+1)+I(I+1)]/S(S+1)$ is the nuclear slowing down factor with $S=1/2$ and $I$ the nuclear spin [52]. Using the SD cross sections listed in [52] and the respective relative velocities $\overline{v}$ we find for a number density $n=10^{14}~{}{\rm cm^{-3}}$ and volume $V=10~{}{\rm cm^{3}}$ that $\delta B=2\times 10^{-17}~{}{\rm T}$ for ${}^{41}{\rm K}$, $\delta B=10^{-16}~{}{\rm T}$ for ${}^{87}{\rm Rb}$, and $\delta B=10^{-14}~{}{\rm T}$ for ${}^{133}{\rm Cs}$. Using instead the parameters of [57], where the authors used a cesium vapor of volume $V=1~{}{\rm cm^{3}}$ and number density $2\times 10^{13}~{}{\rm cm^{-3}}$, we find again $\delta B\approx 10^{-14}~{}{\rm T}$. Since this noise is distributed within the bandwidth $1/\tau$, the corresponding spectral density is $10~{}{\rm pG/\sqrt{Hz}}$, whereas the authors measured a noise level of $400~{}{\rm pG/\sqrt{Hz}}$, and under the premise of a theoretical optimization project a noise level of $2~{}{\rm pG/\sqrt{Hz}}$. Finally, using (3) we can obtain the ERL which reads
@@
There is yet another interpretation for the factor $\kappa$ in (5), interesting in its own right. We can consider the vacuum permeability $\mu_{0}$ is replaced by $\kappa\mu_{0}$, as if the spins have a tendency to align, and $\kappa$ represents the relative permeability constant. But indeed, binary spin-dependent collisions do have the tendency to correlate atoms, as was shown in [53] for the case of spin-exchange collisions. As mentioned before, spin-destruction collisions are described by a spin phase change $\phi$, which is $\phi\ll 1$. In [53] it was shown that the negativity of the two-atom spin state scales as $\sin\phi_{\rm se}^{2}$, where $\phi_{\rm se}\approx 1$ is the spin-exchange phase change. Similarly, here we can state that the entanglement between two colliding atoms will scale like $\sin\phi^{2}\approx\phi^{2}$. Hence any two colliding partners will share a small correlation. The upside is that this correlation can be shared by many more atoms. Imagine a ”bath” atom experiencing consecutive collisions with many ”system” atoms. These will all be slightly correlated, and will reside in a ”correlation volume” $V_{c}$. Consider two spins inside this volume, which are about to collide. Their interaction energy is $\epsilon=\kappa\mu_{0}\mu^{2}/\tilde{V}$, where $\tilde{V}$ is the volume defined by the two spins. This interaction reorients the spins in a time scale $\hbar/\epsilon$. For the aforementioned correlation to be maintained, this time should be equal or larger than $\tilde{V}/\sigma_{\rm sd}\overline{v}$, which is the time required for their collision to correlate them. From this requirement we obtain again $\kappa=\hbar\sigma_{\rm sd}\overline{v}/\mu_{0}\mu^{2}$. We can also estimate the correlation volume $V_{c}$, which will contain $N_{c}$ atoms. We can write $\kappa=\sqrt{N_{c}}\phi$, since the spin variance scales as $\phi^{2}$ and thus the uncertainty as $\phi$, while for $\phi\approx 1$ we would get an enhancement $\kappa=\sqrt{N_{c}}$, retrieving the case of fully polarized spins. The phase $\phi$ is related to the SD relaxation time and the collision rate $1/T$, i.e. it is $1/\tau=\phi^{2}/T$ as discussed in [56]. The collision rate is $1/T=\overline{v}/n^{-1/3}$. Putting everything together we find $N_{c}=(\hbar\overline{v}/\mu_{0}\mu^{2})^{2}\sigma_{\rm sd}n^{-2/3}$. For a ${}^{41}{\rm K}$ number density of $n=10^{14}~{}{\rm cm}^{-3}$ it is $N_{c}\approx 10^{9}$, and $V_{c}=N_{c}/n\approx 0.01~{}{\rm mm}^{3}$. For ${}^{133}{\rm Cs}$ these numbers would be $N_{c}\approx 10^{13}$, and $V_{c}\approx 0.1~{}{\rm cm}^{3}$.
@@
The physical picture that emerges is that both a polarized vapor being relaxed, as well as an unpolarized vapor, will behave like a “squashy” spin medium exhibiting spin fluctuations, split in a number of “domains” not unlike ferromagnetic systems. Such domains exhibit intra-domain correlations, but not inter-domain correlation. The vapor as a whole continuously exchanges energy with its self-field, setting an unavoidable noise level for measuring externally applied magnetic fields. Towards an experimental verification of the noise $\delta B$, we propose the use of a miniaturized cesium cell of volume e.g. $1~{}{\rm mm^{3}}$, in order to boost the previous estimates of $\delta B$ by a factor of $10^{3}$. The cell should be surrounded by a superconducting flux transformer in order to alleviate Johnson noise that would dominate the signal induced by the changing flux produced by $\delta B$ in an ordinary coil.